OpenStudy (anonymous):

Need Help will award Medal! 5x - y = 8 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points)

OpenStudy (clalgee):

I'll try to help you :)

OpenStudy (anonymous):

TY!

OpenStudy (clalgee):

Alright, so for Part A: You can convert it to make it look more like a linear equation. Example: y=mx+b. The number of solutions is infinite because there is no limit to what you can put for x to get y. Does that make sense?

OpenStudy (anonymous):

yes i understand A. i just really need help with B and C

OpenStudy (clalgee):

For Part B: You have to plug in three different numbers for x to get y. You have to write all three in this form (x,y). So I THINK it may be this- (3,7)----5(3)-7=8.

OpenStudy (clalgee):

For Part C: Inconsistent equations R parallel lines. So from find the slope from A and make up your own with the same slope. (3,7)----5(3)-7=8. For Part B: (2,4)-----5(2)-4=8 (My mistake! Sorry!)

OpenStudy (anonymous):

how did you come up with (2,4)---5(2)-4=8? and (3,7)---5(3)-7=8?

OpenStudy (anonymous):

im a little slow, i know :/

OpenStudy (texaschic101):

5x - y = 8 just sub any number in for x and find y. make x = 1 5(1) - y = 8 5 - y = 8 5 - 8 = y - 3 = y so when x = 1, y = -3 now pick another number to plug on for x x = 2 5(2) - y = 8 10 - y = 8 10 - 8 = y 2 = y so when x = 2, then y = 2 now you can pick a number for x and solve for y understand ?

OpenStudy (texaschic101):

as for C....for them to be inconsistent, they cannot intersect...they are parallel. Parallel lines will have the same slope, but different y intercepts.

OpenStudy (anonymous):

x=3 5(3)-y=8 15-y=8 15-8=y y=7 x=3,y=7 (:

OpenStudy (anonymous):

how can i make them not intersect?

OpenStudy (texaschic101):

x = 3, y = 7 ....correct

OpenStudy (texaschic101):

parallel lines will have same slope and different y intercepts